Find the potential when F =-Vº, a conservative force of field defined by F = (3x’yz – 3yli + (xºz – 3x)j + (x*y+2z)k
Answer #1

# solution it E = (38°42-3y) i + (x37-38); + (x24 +22] K A But ginen F = -10 fixa + id q + ů a sy ax 27 F a ax 29 x pi î -208177) = -27 g(z) 2 Jzdz z dz = -2.72 + c 2 y constant g(2) = -72+C therefore from @ LE z²+ c 3xy - x 3 y z potential Holbroal